# Towards Sign Distance Function based Metamaterial Design: Neural Operator Transformer for Forward Prediction and Diffusion Models for Inverse Design

### v1

@@ Abstract
The inverse design of metamaterial architectures presents a significant challenge, particularly for nonlinear mechanical properties involving large deformations, buckling, contact, and plasticity. Traditional methods, such as gradient-based optimization, and recent generative deep-learning approaches often rely on binary pixel-based representations, which introduce jagged edges that hinder finite element (FE) simulations and 3D printing. To overcome these challenges, we propose an inverse design framework that utilizes a signed distance function (SDF) representation combined with a conditional diffusion model. The SDF provides a smooth boundary representation, eliminating the need for post-processing and ensuring compatibility with FE simulations and manufacturing methods. A classifier-free guided diffusion model is trained to generate SDFs conditioned on target macroscopic stress-strain curves, enabling efficient one-shot design synthesis. To assess the mechanical response of the generated designs, we introduce a forward prediction model based on Neural Operator Transformers (NOT), which accurately predicts homogenized stress-strain curves and local solution fields for arbitrary geometries with irregular query meshes. This approach enables a closed-loop process for general metamaterial design, offering a pathway for the development of advanced functional materials.
@@ 2.3 Inverse design by diffusion model
In this section, we describe the classifier-free guided diffusion model [40] trained to design the micro-structure of the unit cell of the metamaterial that can achieve a target stress-strain curve. The label data $y$ is the stress-strain curve obtained from the FEM simulation of the corresponding geometry, represented by SDF ($x_{0}$).
@@
While our developed diffusion model performs well on the test dataset for inverse retrieval, real-world applications often demand on-the-fly inverse design of custom-defined material properties with high fidelity. Therefore, instead of using stress-strain curves from FEM simulations as shown in Extended Data Fig. 2, a robust inverse design model should be able to generate possible designs from custom and predefined stress-strain curves. Thus, we now use the Ramberg-Osgood equation [41] to define the target stress-strain curve. The Ramberg-Osgood equation is defined as
@@
Using the Ramberg-Osgood equations, we generate two target stress-strain curves with different material properties, as shown in Fig. 5(A) and (B). Both targets use the same hardening exponent $n=10$ and coefficient $\alpha=0.002$, but differ in Young’s modulus and reference yield stress. Case (A) is obtained using Eq. 7 with a Young’s modulus $E=800$ MPa and yield stress $\sigma_{0}=30$ MPa, while Case (B) is derived from Eq. 8 with $E=1000$ MPa, $\sigma_{0}=40$ MPa, $E^{\prime}=500$ MPa, and $\varepsilon_{c}=0.156$. These target curves are fed into the inverse diffusion model to generate 500 geometry solutions, which are then evaluated using the forward model to predict the stress-strain curves. The four best designs based on the $L_{2}$ relative error are shown in Fig. 5 (c-f). Fig. 5(a) and (b) compare the target and designed stress-strain curves for the four geometries using the forward model and FEM, respectively, with the $L_{2}$ relative error indicated in the curve legend. As shown in Fig. 5(a), the designed geometries closely match the custom-defined stress-strain curves. The FEM simulations of the designed geometries, shown in Fig. 5(b), confirm the accuracy of the designs. Additionally, we predict the solution fields for the designed geometry in Fig. 5(A)(e) at different strain steps, with results shown in Fig. 6. For case (B), the solution fields for the designed geometry (e) are shown in Extended Data Fig. 5. These results demonstrate that the designed geometries achieve the target stress-strain curves and accurately predict the solution fields.
@@ 3 Discussion
The integration of additive manufacturing and advanced computational techniques has opened new avenues for the design and synthesis of metamaterials with tailored mechanical properties with applications in various engineering disciplines including aerospace, automotive, and biomedical. One of the primary challenges addressed in this work is the complexity of inverse design in achieving desired macroscopic mechanical properties, especially in nonlinear scenarios. Traditional methods, such as topology optimization, although effective, are often computationally intensive and prone to local minima due to their reliance on iterative gradient-based optimization schemes. Recently, generative deep learning models have shown potential in inverse design by offering a more efficient approach that eliminates the need for multiple iterations, thus providing a pathway for exploring the vast design space of metamaterials. However, existing methods typically rely on binary pixel-based geometries, which have jagged boundaries and require additional boundary smoothing, and are thus not directly applicable to FEM simulations and 3D printing. In contrast, our proposed method designs architected metamaterials using a classifier-free guided diffusion model, representing geometries through signed distance functions (SDFs). This approach enables the generation of multiple SDF solutions for a given target stress-strain curve. The transition from binary pixel-based representations to an SDF framework marks a significant advancement, as SDFs provide smooth boundary representations that eliminate jagged edges. This refinement reduces mesh distortions in finite element simulations and simplifies the 3D printing process, ultimately improving simulation accuracy, convergence, and manufacturability, and ensuring the robustness and reliability of the designed structures.
@@ 4 Methods
This section provides details on the generation of random unit cell geometries, FEM simulations, and the implementation of deep learning models. Additional explanations can be found in the Supplementary Information.
@@ 8 Author contributions statement
Q. Liu conceived the concept, generated the dataset, developed the methodology, implemented the models, performed the analysis, and drafted, reviewed, and edited the manuscript. S. Koric contributed to the results discussion and to the configuration of Abaqus on HPC, and reviewed and edited the manuscript. D. Abueidda and H. Meidani contributed to the result discussion and reviewed and edited the manuscript. P. Geubelle conceived the concept, supervised the project, contributed to the result discussion, and reviewed and edited the manuscript.
@@ 9 Competing interests statement
S1 Binary pixel-based vs SDF-based geometry representation
@@
$U\left(x\right)=\sum_{i=1}^{N}\sqrt{2S(k_{i})\Delta k}\left(Z_{1,i}\cdot\cos\left(\left\langle k_{i},x\right\rangle\right)+Z_{2,i}\cdot\sin\left(\left\langle k_{i},x\right\rangle\right)\right)$ (S1)
@@
For each generated unit cell, we store the vertices’ coordinates and the contour connectivities (outer boundary and holes) in a json file, which can be read by Abaqus using the Python API for building geometry and meshing. To evaluate the SDF for each unit cell, we first build a polygon object using shapely.geometry from the open-source Python package shapely, then calculate the shortest distance of $120\times 120$ uniform grid points in $[-0.1~{}\text{mm},1.1~{}\text{mm}]^{2}$ to the polygon.
@@
The domain material is acetal homopolymer resin [3], and an elastic-plastic constitutive model is used to describe the material behavior. The Young’s modulus of the resin is 2.3 GPa, and the Poisson’s ratio is 0.35. The plastic behavior is described by a piece-wise linear hardening model as shown in Table S1, and the corresponding stress-strain curve is shown in Fig. S2.
@@
Figure S2: stress-strain curve for homopolymer resin
@@
We conducted a mesh sensitivity analysis to determine the optimal mesh size for the finite element simulation. The mesh size around the boundary contour is approximately 1/64 mm if the global mesh size is larger than that, as the geometry contour is extracted from a GRF with a resolution of 64x64. We found that a mesh size of 0.04 mm is sufficient to capture the material response of the unit cell. The stress-strain curve obtained using a mesh size of 0.04 mm is compared with the curve obtained using mesh sizes of 0.01 mm, 0.02 mm, and 0.08 mm as shown in Fig. S3. The results indicate that the stress-strain curves obtained using the four mesh sizes are almost identical, confirming that a mesh size of 0.04 mm is adequate to capture the material response of the unit cell.
Figure S3: Mesh sensitivity analysis for one unit cell
@@
Based on the random, arbitrary periodic unit cell generation and the corresponding FEM simulation, we generate 10K samples. Training the forward neural networks exhibits severe overfitting due to the limited number of training samples and the complexity of the data. To alleviate this issue, we apply data augmentation to the dataset. For data augmentation, we randomly shift the unit cell in the x direction multiple times. The new unit cell $[x_{0},x_{0}+1]\times[0,1]~{}\text{mm}^{2}$ is created by shifting the original unit cell by $x_{0}$, where $x_{0}$ is randomly selected from $[0.1,0.9]$ mm. The solution field (stress, displacement) of the new unit cell is the same as the original one but with shifted node coordinates by $x_{0}$:
$x_{\text{new}}=\begin{cases}x_{\text{original}}+1-x_{0}&\text{if }x_{\text{original}}<x_{0}\\ x_{\text{original}}-x_{0}&\text{if }x_{\text{original}}\geq x_{0}\end{cases}$ (S2)
@@
Table S2: Geometry encoder architecture of the forward NOT model
@@ 26 Reshape + Permute (Output) (B,225,64)
Table S3: Solution decoder architecture of the forward NOT model. $N$ represents the number of query points. For stress-strain curve prediction, $N=51$, and for solution fields prediction, it represents the maximum number of nodes of samples in the batch. $C_{in}$ represents the input channels, and is equal to 1 for stress-strain curve prediction, and 2 for solution fields prediction. $C_{out}$ represents the output channels, and is equal to 1 for stress-strain curve prediction, and $3\times 26$ for solution fields (Mises stress and displacements at 26 strain steps) prediction.
@@ 6 Linear (Output) (B, N, $C_{out}$)
The mean square error (MSE) is used as the loss function for training the forward NOT models. The loss function is defined as:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (S3)
When training the forward NOT model for predicting the solution fields, because the number of query points is padded to the maximum number of nodes in the batch, we apply a mask mechanism to exclude the padding points from the loss calculation:
@@
where $m_{i}$ is the mask value, which is 1 if the point is not padding, and 0 if the point is padding.
The training hyperparameters for the forward NOT models are summarized in Table S4.
Table S4: Training hyperparameters of the forward NOT models
@@
The performance of the forward NOT model for predicting the solution fields is shown in Fig. S5. In Fig. S5(a), the MSE loss during training is presented, while Fig. S5(b) displays the $L_{2}$ relative error distribution for the test data. The mean $L_{2}$ relative error is 10.3% with a standard deviation of 4.6%, indicating that the model generalizes well to the test data. The best and worst prediction cases from the test data are illustrated in Fig. S6 and Fig. S7, respectively. The $L_{2}$ relative error over the total 26 strain steps is 3.9% for the best case and 72.6% for the worst case.
@@
In this work, the classifier-free guided diffusion model [5] is trained to design the micro-structure of the unit cell of a metamaterial that can achieve a target stress-strain curve. The labeled data $y$ consists of the stress-strain curve obtained from the ABAQUS simulation of the corresponding random geometry. The corresponding SDF image serves as the design variable $x_{0}$ for the diffusion model. By applying the marching algorithm on the designed SDF images to extract the contour with a value equal to 0, the boundary of the designed geometry is obtained. The model architecture is shown in Fig. 2(d). For completeness, the denoising diffusion probabilistic model [4] and the classifier-free guidance method [5] are briefly reviewed here.
@@
$p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)\right),$ (S8)
The reverse denoising process is then controlled by,
@@
The optimization objective of the denoising diffusion probabilistic model is to maximize the log-likelihood $\log p_{\theta}\left(\mathbf{x}_{0:T}\right)$, which results in the following loss function,
@@
Training a NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ and minimizing the loss function $L$ leads to $\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)\approx\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)$, thus making the proposed posterior distribution $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$ close to the true posterior distribution $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$. The above reverse generative process is random and not controlled by any specific target. In specific design tasks, we aim to generate fields that represent the target, such as the force-displacement curve, which requires training the NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ with conditional information. To incorporate the condition information $\mathbf{y}$ into the diffusion process, Ho and Salimans [40] proposed a classifier-free guidance method, incorporating the scores from a conditional diffusion model $p_{\theta}(\mathbf{x}|\mathbf{c})$ and an unconditional diffusion model $p_{\theta}(\mathbf{x})$. The noise estimators $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ of $p_{\theta}(\mathbf{x})$ and $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ of $p_{\theta}(\mathbf{x}|\mathbf{c})$ are trained in a single NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)$. Here,$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ is trained with paired data $(\mathbf{x},\mathbf{c}=\mathbf{y})$, and $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ is trained with data $\mathbf{x}$ only, i.e.,$(\mathbf{x},\mathbf{c}=\emptyset)$. During the training process, the condition information $\mathbf{c}$ is randomly set as $\mathbf{c}=\emptyset$ or $\mathbf{c}=\mathbf{y}$ for sample $(\mathbf{x},\mathbf{y})$ at different time steps $t$, allowing the model to learn to generate fields both conditionally and unconditionally. During the reverse inference process, $\epsilon_{\theta}$ in Eq. S10 is replaced by the linear summation of conditional and unconditional noise estimators,

### v2

# Toward Signed Distance Function based Metamaterial Design: Neural Operator Transformer for Forward Prediction and Diffusion Model for Inverse Design

@@ Abstract
The inverse design of metamaterial architectures presents a significant challenge, particularly for nonlinear mechanical properties involving large deformations, buckling, contact, and plasticity. Traditional methods, such as gradient-based optimization, and recent generative deep-learning approaches often rely on binary pixel-based representations, which introduce jagged edges that hinder finite element (FE) simulations and 3D printing. To overcome these challenges, we propose an inverse design framework that utilizes a signed distance function (SDF) representation combined with a conditional diffusion model. The SDF provides a smooth boundary representation, eliminating the need for post-processing and ensuring compatibility with FE simulations and manufacturing methods. A classifier-free guided diffusion model is trained to generate SDFs conditioned on target macroscopic stress-strain curves, enabling efficient one-shot design synthesis. To assess the mechanical response and the quality of the generated designs, we introduce a forward prediction model based on Neural Operator Transformers (NOT), which accurately predicts homogenized stress-strain curves and local solution fields for arbitrary geometries with irregular query meshes. This approach enables a closed-loop process for general metamaterial design, offering a pathway for the development of advanced functional materials.
@@ 2.3 Inverse design by diffusion model
In this section, we describe the classifier-free guided diffusion model [40] trained to design the micro-structure of the unit cell of the metamaterial that can achieve a target stress-strain curve. The label data $y$ is the stress-strain curve obtained from the FE simulation of the corresponding geometry, represented by SDF ($x_{0}$).
@@
While our developed diffusion model performs well on the test dataset for inverse retrieval, real-world applications often demand on-the-fly inverse design of custom-defined material properties with high fidelity. Therefore, instead of using stress-strain curves from FE simulations as shown in Extended Data Fig. 2, a robust inverse design model should be able to generate possible designs from custom and predefined stress-strain curves. Thus, we now use the Ramberg-Osgood equation [41] to define the target stress-strain curve. The Ramberg-Osgood equation is defined as
@@
Using the Ramberg-Osgood equations, we generate two target stress-strain curves with different material properties, as shown in Fig. 5(A) and (B). Both targets use the same hardening exponent $n=10$ and coefficient $\alpha=0.002$, but differ in Young’s modulus and reference yield stress. Case (A) is obtained using Eq. 7 with a Young’s modulus $E=800$ MPa and yield stress $\sigma_{0}=30$ MPa, while Case (B) is derived from Eq. 8 with $E=1000$ MPa, $\sigma_{0}=40$ MPa, $E^{\prime}=500$ MPa, and $\varepsilon_{c}=0.156$. These target curves are fed into the inverse diffusion model to generate 500 geometry solutions, which are then evaluated using the forward model to predict the stress-strain curves. The four best designs based on the $L_{2}$ relative error are shown in Fig. 5 (c-f). Fig. 5(a) and (b) compare the target and designed stress-strain curves for the four geometries using the forward model and FE simulations, respectively, with the $L_{2}$ relative error indicated in the curve legend. As shown in Fig. 5(a), the designed geometries closely match the custom-defined stress-strain curves. The FE simulations of the designed geometries, shown in Fig. 5(b), confirm the accuracy of the designs. Additionally, we predict the solution fields for the designed geometry in Fig. 5(A)(e) at different strain steps, with results shown in Fig. 6. For case (B), the solution fields for the designed geometry (e) are shown in Extended Data Fig. 5. These results demonstrate that the designed geometries achieve the target stress-strain curves and accurately predict the solution fields.
@@ 3 Discussion
The integration of additive manufacturing and advanced computational techniques has opened new avenues for the design and synthesis of metamaterials with tailored mechanical properties with applications in various engineering disciplines including aerospace, automotive, and biomedical. One of the primary challenges addressed in this work is the complexity of inverse design in achieving desired macroscopic mechanical properties, especially in nonlinear scenarios. Traditional methods, such as topology optimization, although effective, are often computationally intensive and prone to local minima due to their reliance on iterative gradient-based optimization schemes. Recently, generative deep learning models have shown potential in inverse design by offering a more efficient approach that eliminates the need for multiple iterations, thus providing a pathway for exploring the vast design space of metamaterials. However, existing methods typically rely on binary pixel-based geometries, which have jagged boundaries and require additional boundary smoothing, and are thus not directly applicable to FE simulations and 3D printing. In contrast, our proposed method designs architected metamaterials using a classifier-free guided diffusion model, representing geometries through signed distance functions (SDFs). This approach enables the generation of multiple SDF solutions for a given target stress-strain curve. The transition from binary pixel-based representations to an SDF framework marks a significant advancement, as SDFs provide smooth boundary representations that eliminate jagged edges. This refinement reduces mesh distortions in finite element simulations and simplifies the 3D printing process, ultimately improving simulation accuracy, convergence, and manufacturability, and ensuring the robustness and reliability of the designed structures.
@@ 4 Methods
This section provides details on the generation of random unit cell geometries, FE simulations, and the implementation of deep learning models. Additional explanations can be found in the Supplementary Information.
@@ 8 Author contributions statement
Q. Liu conceived the concept, generated the dataset, developed the methodology, implemented the models, performed the analysis, and drafted, reviewed, and edited the manuscript. S. Koric contributed to the results discussion and to the configuration of Abaqus on HPC, and reviewed and edited the manuscript. D. Abueidda and H. Meidani contributed to the result discussion and reviewed and edited the manuscript. P. Geubelle proposed the concept, supervised the project, contributed to the result discussion, and reviewed and edited the manuscript.
@@ 9 Competing interests statement
S1 Binary pixel-based vs. SDF-based geometry representation
@@
$U\left(x\right)=\sum_{i=1}^{N}\sqrt{2S(k_{i})\Delta k}\left(Z_{1,i}\cdot\cos\left(\left\langle k_{i},x\right\rangle\right)+Z_{2,i}\cdot\sin\left(\left\langle k_{i},x\right\rangle\right)\right),$ (S1)
@@
For each generated unit cell, we store the coordinates of the vertices and the contour connectivities (outer boundary and holes) in a json file, which can be read by Abaqus using the Python API for building geometry and meshing. To evaluate the SDF for each unit cell, we first build a polygon object using shapely.geometry from the open-source Python package shapely, then calculate the shortest distance of $120\times 120$ uniform grid points in $[-0.1~{}\text{mm},1.1~{}\text{mm}]^{2}$ to the polygon.
@@
The material adopted in this work is acetal homopolymer resin [3] and an elastic-plastic model is used to describe its constitutive behavior. The Young’s modulus of the resin is 2.3 GPa and the Poisson’s ratio is 0.35. The plastic behavior is described by a piecewise linear hardening model presented in Table S1 and the corresponding stress-strain curve is shown in Fig. S2.
@@
Figure S2: Stress-strain curve for homopolymer resin.
@@
We conducted a mesh sensitivity analysis to determine the optimal mesh size for the finite element simulations. It is important to note that the domain boundary consists of small segments, each approximately 1/64 mm in length. This segment length corresponds to the resolution of the Gaussian random field (GRF) used to generate the geometry. Consequently, the mesh size around the contour does not exceed the segment length. Our analysis indicates that a global mesh size of 0.04 mm is sufficient to capture the compression response of the unit cells. The stress-strain curves obtained using a mesh size of 0.04 mm are compared with the curves obtained using mesh sizes of 0.01 mm, 0.02 mm, and 0.08 mm as shown in Fig. S3. The results indicate that the stress-strain curves obtained using the four mesh sizes are almost identical, confirming that a mesh size of 0.04 mm is adequate to capture the material response of the unit cell.
Figure S3: Mesh sensitivity analysis for two unit cells under compression: (a) without self-contact and (b) with self-contact.
@@
Based on the random, arbitrary periodic unit cell generation and the corresponding FE simulation, we generate 10K samples. Training the forward neural networks exhibits severe overfitting due to the limited number of training samples and the complexity of the data. To alleviate this issue, we apply data augmentation to the dataset by randomly shifting the unit cell in the $x$-direction multiple times. The new unit cell $[x_{0},x_{0}+1]\times[0,1]~{}\text{mm}^{2}$ is created by shifting the original unit cell horizontally by $x_{0}$, where $x_{0}$ is randomly selected from $[0.1,0.9]$ mm. The solution stress and displacement fields of the new unit cell are the same as the original one but with the node coordinates shifted by $x_{0}$:
$x_{\text{new}}=\begin{cases}x_{\text{original}}+1-x_{0}&\text{if }x_{\text{original}}<x_{0},\\ x_{\text{original}}-x_{0}&\text{if }x_{\text{original}}\geq x_{0}.\end{cases}$ (S2)
@@
Table S2: Geometry encoder architecture of the forward NOT model.
@@ 27 Reshape + Permute (Output) (B,225,64)
Table S3: Solution decoder architecture of the forward NOT model. $N$ represents the number of query points. For the stress-strain curve prediction, $N=51$, while for solution fields prediction, $N$ represents the maximum number of sampled nodes of in the batch. $C_{in}$ represents the input channels and is equal to 1 for stress-strain curve prediction and 2 for solution fields prediction. $C_{out}$ corresponds to the output channels and is equal to 1 for stress-strain curve prediction, and $3\times 26$ for solution field prediction (Mises stress and displacements at 26 strain steps).
@@ 7 Linear (Output) (B, N, $C_{out}$)
The mean square error (MSE) is used as the loss function for training the forward NOT models. The loss function is defined as
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}-\hat{y}_{i}\right)^{2}.$ (S3)
When training the forward NOT model for predicting the solution fields, as the number of query points is padded to the maximum number of nodes in the batch, we apply a mask mechanism to exclude the padding points from the loss calculation:
@@
where $m_{i}$ is the mask value, which is 1 if it is not a padded point, and 0 if it is. The training hyperparameters for the forward NOT models are summarized in Table S4.
Table S4: Training hyperparameters of the forward NOT models.
@@
The performance of the forward NOT model for predicting the solution fields is presented in Fig. S5. Fig. S5(a) shows the MSE loss during training, while Fig. S5(b) displays the $L_{2}$ relative error distribution for the test data. The mean $L_{2}$ relative error is 10.3% with a standard deviation of 4.6%, indicating that the model generalizes well to the test data. The best and worst prediction cases from the test data are illustrated in Fig. S6 and Fig. S7, respectively. The $L_{2}$ relative error over the total 26 strain steps is 3.9% for the best case and 72.6% for the worst case.
@@
In this work, the classifier-free guided diffusion model [5] is trained to design the micro-structure of the unit cell of a metamaterial that can achieve a target stress-strain curve. The labeled data $y$ consists of the stress-strain curve obtained from the ABAQUS simulation of the corresponding random geometry. The corresponding SDF image serves as the design variable $\mathbf{x_{0}}$ for the diffusion model. By applying the marching algorithm on the designed SDF images to extract the contour with a value equal to 0, the boundary of the designed geometry is obtained. The model architecture is shown in Fig. 2(d). For completeness, the denoising diffusion probabilistic model [4] and the classifier-free guidance method [5] are briefly reviewed here.
@@
$p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right),\mathbf{\Sigma}_{\theta}\left(\mathbf{x}_{t},t\right)\right).$ (S8)
The reverse denoising process is then controlled by
@@
The optimization objective of the denoising diffusion probabilistic model is to maximize the log-likelihood $\log p_{\theta}\left(\mathbf{x}_{0:T}\right)$, which results in the following loss function:
@@
Training a NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ and minimizing the loss function $L$ leads to $\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)\approx\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)$, thus making the proposed posterior distribution $p_{\theta}\left(\mathbf{x}_{t-1}|\mathbf{x}_{t}\right)$ close to the true posterior distribution $q\left(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0}\right)$. The above reverse generative process is random and not controlled by any specific target. In specific design tasks, we aim to generate fields that represent the target, such as the force-displacement curve, which requires training the NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ with conditional information. To incorporate the condition information $\mathbf{y}$ into the diffusion process, Ho and Salimans [40] proposed a classifier-free guidance method that incorporates the scores from a conditional diffusion model $p_{\theta}(\mathbf{x}|\mathbf{c})$ and an unconditional diffusion model $p_{\theta}(\mathbf{x})$. The noise estimators $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ of $p_{\theta}(\mathbf{x})$ and $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ of $p_{\theta}(\mathbf{x}|\mathbf{c})$ are trained in a single NN $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}\right)$. Here, $\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{c}=\mathbf{y}\right)$ is trained with paired data $(\mathbf{x},\mathbf{c}=\mathbf{y})$, and $\epsilon_{\theta}\left(\mathbf{x}_{t},t\right)$ is trained with data $\mathbf{x}$ only, i.e.,$(\mathbf{x},\mathbf{c}=\emptyset)$. During the training process, the condition information $\mathbf{c}$ is randomly set as $\mathbf{c}=\emptyset$ or $\mathbf{c}=\mathbf{y}$ for sample $(\mathbf{x},\mathbf{y})$ at different time steps $t$, allowing the model to learn to generate fields both conditionally and unconditionally. During the reverse inference process, $\epsilon_{\theta}$ in Eq. S10 is replaced by the linear summation of conditional and unconditional noise estimators,